# Chiara Amorino, Vytautė Pilipauskaitė

### v2

@@ 5 Preparation for the proofs
We now state a bound on the error committed moving from $X^{i}_{j\Delta_{n}}$ to $\tilde{X}^{i}_{j\Delta_{n}}$ and approximating $\tilde{X}^{i}_{j\Delta_{n}}-\tilde{X}^{i}_{(j-1)\Delta_{n}}$ by $\Delta_{n}b(\tilde{Z}^{i}_{j\Delta_{n}},\tilde{\Pi}^{N}_{j\Delta_{n}})$. Its proof is in Section 7. For $i=1,\dots,N$, $j=0,\dots,n-1$, let us introduce
@@
Additionally, it is important to note that this $\Delta_{n}^{\frac{1}{2}}$ correlation arises from a stochastic integral, which can be approximated using the Euler-Maruyama scheme, as seen in the right hand side of (i). Crucially, this correlation term is conditionally centered, which plays a significant role in our analysis. It allows us to consider the following term in the development in (i), consisting in $\tilde{\varepsilon}_{j}^{i}$, which is now of order $\Delta_{n}$.
@@
The results for complete observations extend to partial observations thanks to condition (4). The approximation of $\tilde{X}^{i}_{(j+1)\Delta_{n}}-\tilde{X}^{i}_{j\Delta_{n}}$ in Proposition 5.4(i) is also useful in the proofs of Theorems 5.9 and 5.10, for partial observations. Complete observations satisfy a similar approximation (see Lemma 5.3 and its proof in [4]), but with $V^{i}_{j}:=\Delta_{n}^{-1/2}(B^{i}_{(j+1)\Delta_{n}}-B^{i}_{j\Delta_{n}})$ replacing the error factor $U^{i}_{j}$. We note that $\mathbb{E}[(U^{i}_{j})^{2}]=2/3$, whereas $\mathbb{E}[(V^{i}_{j})^{2}]=1$. Therefore, we have an additional factor of $2/3$ in Theorem 5.10 for partial observations.
The convergence in probability established in the theorems above is essential for proving the consistency of our estimators. We now turn to preparing the proof of asymptotic normality, which relies on the following convergence in distribution. We begin by analyzing $I_{n}^{N,C}(f_{\theta})$ and $I_{n}^{N,P}(f_{\theta})$.
@@
$\frac{1}{\sqrt{N}}I_{n}^{N,C}(f)\xrightarrow{d}N(0,\bar{\Pi}((fa_{\sigma_{0}})^{2})),\qquad\frac{1}{\sqrt{N}}I_{n}^{N,P}(f)\xrightarrow{d}N(0,\bar{\Pi}((fa_{\sigma_{0}})^{2})).$
@@
$\frac{1}{\sqrt{N\Delta_{n}}}(Q_{n}^{N,C}(f)-\Delta_{n}\nu_{n}^{N,C}(fa_{\sigma_{0}}^{2}))\xrightarrow{d}N(0,2\bar{\Pi}(f^{2}a^{4}_{\sigma_{0}})),$
$\frac{1}{\sqrt{N\Delta_{n}}}(Q_{n}^{N,P}(f)-\frac{2}{3}\Delta_{n}\nu_{n}^{N,P}(fa_{\sigma_{0}}^{2}))\xrightarrow{d}N(0,\bar{\Pi}(f^{2}a^{4}_{\sigma_{0}})).$
@@
6 Proof main results
@@
$\frac{\Delta}{N}L_{n}^{N,P}(\mu,\sigma)$ $\xrightarrow{\mathbb{P}}\bar{\Pi}\Big{(}\frac{a^{2}_{\sigma_{0}}}{a^{2}_{\sigma}}+\log a^{2}_{\sigma}\Big{)}.$ (19)
@@
uniformly in $\theta$ according to Proposition 5.8 and Theorems 5.9 and 5.10. Similarly, in the partial observation case we have that
@@
uniformly in $(\mu,\sigma)$ according to the same Proposition 5.8 and Theorems 5.9 and 5.10.
@@
Observe that, in the following, we will provide the proof of our main results in the case where the parameters are mono-dimensional, in order to lighten the notation. The scheme of the proof is the same in both the complete and the partial observation case. For $A\in\{P,C\}$, we use Taylor’s formula and, as $\nabla_{\theta}L_{n}^{N,A}(\hat{\theta}_{n}^{N,A})=0$, we obtain
@@
$\Sigma_{n}^{N,A}(\theta)$ $=M_{n}^{N,A}\nabla^{2}_{\theta}L_{n}^{N,A}(\theta_{0})M_{n}^{N,A}$
@@
Note that, for simplicity, we consider the parameters $\mu$ and $\sigma$ in one dimension throughout this discussion to streamline the notation. However, the analysis naturally extends to the multidimensional case. The asymptotic behaviour of the estimator is given by the asymptotic behaviour of the first and second derivatives of the contrast function. Let us now detail the cases of complete and partial observation separately. We start considering the complete observation framework. Note that
@@
$-\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,C}(\theta_{0})$ $=\frac{1}{\sqrt{N\Delta}}Q_{n}^{N,C}\Big{(}\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{4}_{\sigma_{0}}}\Big{)}-2\sqrt{\frac{\Delta}{N}}I_{n}^{N,C}\Big{(}\frac{b_{\mu_{0}}\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{4}_{\sigma_{0}}}\Big{)}-\sqrt{\frac{\Delta}{N}}\nu_{n}^{N,C}\Big{(}\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{2}_{\sigma_{0}}}\Big{)}$
$\qquad-\Delta\sqrt{\frac{\Delta}{N}}\nu_{n}^{N,C}\Big{(}\frac{\partial_{\sigma}c_{\sigma_{0}}b_{\mu_{0}}^{2}}{c^{2}_{\sigma_{0}}}\Big{)}$
$=\frac{1}{\sqrt{N\Delta}}Q_{n}^{N,C}\Big{(}\frac{\partial_{\sigma}c_{\sigma_{0}}}{c^{2}_{\sigma_{0}}}\Big{)}-\sqrt{\frac{\Delta}{N}}\nu_{n}^{N,C}\Big{(}\frac{\partial_{\sigma}c_{\sigma_{0}}}{c_{\sigma_{0}}}\Big{)}+o_{\mathbb{P}}(1).$
Then, Theorem 5.11 ensures that
$\frac{1}{\sqrt{N}}\partial_{\mu}L_{n}^{N,C}(\theta_{0})\xrightarrow{d}N(0,\bar{\Pi}((2\frac{\partial_{\mu}b_{\mu_{0}}}{c_{\sigma_{0}}}a_{\sigma_{0}})^{2}))=N(0,4\bar{\Pi}((\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}})^{2})).$
@@
$\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,C}(\theta_{0})\xrightarrow{d}N(0,2\bar{\Pi}((\frac{\partial_{\sigma}c_{\sigma_{0}}}{c_{\sigma_{0}}})^{2})).$
@@
$\nabla_{\theta}L_{n}^{N,C}(\theta_{0})M_{n}^{N,C}\xrightarrow{d}N(0,\Sigma^{C}(\theta_{0})),$
@@
$\Sigma^{C}(\theta_{0})=\begin{pmatrix}4\bar{\Pi}((\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}})^{2})&0\\ 0&2\bar{\Pi}((\frac{\partial_{\sigma}c_{\sigma_{0}}}{c_{\sigma_{0}}})^{2})\end{pmatrix}$.
@@
$\frac{1}{N}\partial^{2}_{\mu}L_{n}^{N,C}(\theta)$ $=-\frac{2}{N}I_{n}^{N,C}(\frac{\partial_{\mu}^{2}b_{\mu}}{c_{\sigma}})+\frac{\Delta}{N}\nu_{n}^{N,C}(\partial_{\mu}(2\frac{\partial_{\mu}b_{\mu}}{c_{\sigma}}(b_{\mu}-b_{\mu_{0}})))$
$=o_{\mathbb{P}}(1)+\frac{\Delta}{N}\nu_{n}^{N,C}(\frac{2}{c_{\sigma}}(\partial_{\mu}^{2}b_{\mu}(b_{\mu}-b_{\mu_{0}})+(\partial_{\mu}b_{\mu})^{2}))$
$\xrightarrow{\mathbb{P}}\bar{\Pi}(\frac{2}{c_{\sigma}}(\partial_{\mu}^{2}b_{\mu}(b_{\mu}-b_{\mu_{0}})+(\partial_{\mu}b_{\mu})^{2}))=:\Sigma_{(1,1)}(\theta)$
uniformly in $\theta$ thanks to Proposition 5.8. Observe in particular that $\Sigma_{(1,1)}(\theta_{0})=\bar{\Pi}(\frac{2(\partial_{\mu}b_{\mu_{0}})^{2}}{c_{\sigma_{0}}})$. Furthermore,
@@
$=2\frac{\sqrt{\Delta}}{N}I_{n}^{N,C}(\frac{\partial_{\mu}b_{\mu}}{(c_{\sigma})^{2}}\partial_{\sigma}c_{\sigma})-\frac{\Delta^{\frac{3}{2}}}{N}\nu_{n}^{N,C}(2\frac{\partial_{\sigma}c_{\sigma}\partial_{\mu}b_{\mu}}{(c_{\sigma})^{2}}(b_{\mu}-b_{\mu_{0}}))\xrightarrow{\mathbb{P}}0$
@@
Regarding the second derivatives in $\sigma$, we have
$\frac{\Delta}{N}\partial^{2}_{\sigma}L_{n}^{N,C}(\theta)=\frac{1}{N}Q_{n}^{N,C}(2\frac{(\partial_{\sigma}c_{\sigma})^{2}}{c_{\sigma}^{3}}-\frac{\partial^{2}_{\sigma}c_{\sigma}}{c^{2}_{\sigma}})+o_{\mathbb{P}}(1)+\frac{\Delta}{N}\nu_{n}^{N,C}(\frac{\partial^{2}_{\sigma}c_{\sigma}}{c_{\sigma}}-\frac{(\partial_{\sigma}c_{\sigma})^{2}}{c_{\sigma}^{2}}),$
that converges in probability to
$\bar{\Pi}(c_{\sigma_{0}}(2\frac{(\partial_{\sigma}c_{\sigma})^{2}}{c_{\sigma}^{3}}-\frac{\partial^{2}_{\sigma}c_{\sigma}}{c^{2}_{\sigma}}))+\bar{\Pi}(\frac{\partial^{2}_{\sigma}c_{\sigma}}{c_{\sigma}}-\frac{(\partial_{\sigma}c_{\sigma})^{2}}{c_{\sigma}^{2}})=:\Sigma_{(2,2)}(\theta)$
uniformly in $\theta$, thanks to Proposition 5.8 and Theorem 5.10. We remark that $\Sigma_{(2,2)}(\theta_{0})=\bar{\Pi}(\frac{(\partial_{\sigma}c_{\sigma_{0}})^{2}}{c_{\sigma_{0}}^{2}})$ This implies that the matrix $\Sigma_{n}^{N,C}(\theta)$ converges in probability, uniformly in $\theta$, towards
$\Sigma(\theta)=\begin{pmatrix}\Sigma_{(1,1)}(\theta)&0\\ 0&\Sigma_{(2,2)}(\theta)\end{pmatrix}$.
@@
$\begin{pmatrix}2\bar{\Pi}(\frac{(\partial_{\mu}b_{\mu_{0}})^{2}}{c_{\sigma_{0}}})&0\\ 0&\bar{\Pi}(\frac{(\partial_{\sigma}c_{\sigma_{0}})^{2}}{c^{2}_{\sigma_{0}}})\end{pmatrix}^{-2}\times\begin{pmatrix}4\bar{\Pi}(\frac{(\partial_{\mu}b_{\mu_{0}})^{2}}{c_{\sigma_{0}}})&0\\ 0&2\bar{\Pi}\frac{(\partial_{\sigma}c_{\sigma_{0}})^{2}}{c^{2}_{\sigma_{0}}}\end{pmatrix}=\begin{pmatrix}\Big{(}\bar{\Pi}(\frac{(\partial_{\mu}b_{\mu_{0}})^{2}}{c_{\sigma_{0}}})\Big{)}^{-1}&0\\ 0&2\Big{(}\bar{\Pi}\frac{(\partial_{\sigma}c_{\sigma_{0}})^{2}}{c^{2}_{\sigma_{0}}}\Big{)}^{-1}\end{pmatrix}$,
@@
$L_{n}^{N,P}(\theta)=\frac{3}{2}\frac{1}{\Delta}Q_{n}^{N,P}(\frac{1}{c_{\sigma}})-3I_{n}^{N,P}(\frac{b_{\mu}}{c_{\sigma}})+\frac{3}{2}\Delta\nu_{n}^{N,P}(\frac{b_{\mu}}{c_{\sigma}}(b_{\mu}-2b_{\mu_{0}}))+\nu_{n}^{N,P}(\log(c_{\sigma})).$
@@
$-\frac{1}{\sqrt{N}}\partial_{\mu}L_{n}^{N,P}(\theta_{0})$ $=\frac{3}{\sqrt{N}}I_{n}^{N,P}(\frac{\partial_{\mu}b_{\mu_{0}}}{c_{\sigma_{0}}})\xrightarrow{d}N(0,9\bar{\Pi}((\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}})^{2})).$
@@
$\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,P}(\theta_{0})\xrightarrow{d}N(0,\frac{9}{4}\bar{\Pi}((\frac{\partial_{\sigma}c_{\sigma_{0}}}{c_{\sigma_{0}}})^{2})).$
Acting as above it is also easy to check that, uniformly in $\theta$,
$\Sigma_{n}^{N,P}(\theta)\xrightarrow{\mathbb{P}}\begin{pmatrix}\frac{3}{2}\Sigma_{11}(\theta)&0\\ 0&\Sigma_{22}(\theta)\end{pmatrix}$
with $\Sigma(\theta)$ defined as in the case of complete observation we dealt with before. We conclude again by applying the inverse of $\int_{0}^{1}\Sigma_{n}^{N,P}(\theta_{0}+u(\hat{\theta}_{n}^{N,P}-\theta_{0}))du$ in Equation (24) and using continuous mapping theorem. The result is therefore proven.
7 Proof technical results

### v3

@@ 6 Preparation for the proofs
We now state a bound on the error committed moving from $X^{i}_{j\Delta_{n}}$ to $\tilde{X}^{i}_{j\Delta_{n}}$ and approximating $\tilde{X}^{i}_{j\Delta_{n}}-\tilde{X}^{i}_{(j-1)\Delta_{n}}$ by $\Delta_{n}b(\tilde{Z}^{i}_{j\Delta_{n}},\tilde{\Pi}^{N}_{j\Delta_{n}})$. Its proof is in Section A. For $i=1,\dots,N$, $j=0,\dots,n-1$, let us introduce
@@
Additionally, it is important to note that this $\Delta_{n}^{\frac{1}{2}}$ correlation arises from a stochastic integral, which can be approximated using the Euler–Maruyama scheme, as seen in the right hand side of (i). Crucially, this correlation term is conditionally centered, which plays a significant role in our analysis. It allows us to consider the following term in the development in (i), consisting in $\tilde{\varepsilon}_{j}^{i}$, which is now of order $\Delta_{n}$.
@@
The results for complete observations extend to partial observations thanks to condition (4). The approximation of $\tilde{X}^{i}_{(j+1)\Delta_{n}}-\tilde{X}^{i}_{j\Delta_{n}}$ in Proposition 6.4(i) is also useful in the proofs of Theorems 6.9 and 6.10, for partial observations. Complete observations satisfy a similar approximation (see Lemma 5.3 and its proof in [4]), but with $V^{i}_{j}:=\Delta_{n}^{-1/2}(B^{i}_{(j+1)\Delta_{n}}-B^{i}_{j\Delta_{n}})$ replacing the error factor $U^{i}_{j}$. We note that $\mathbb{E}[(U^{i}_{j})^{2}]=2/3$, whereas $\mathbb{E}[(V^{i}_{j})^{2}]=1$. Therefore, we have an additional factor of $2/3$ in Theorem 6.10 for partial observations.
The convergence in probability established in the theorems above is essential for proving the consistency of our estimators. It is worth highlighting a direct and highly useful consequence of Theorem 6.10 under partial observations. When the target model features a constant volatility, that is, $a_{\sigma}(x)\equiv\sigma$, one can recover a quadratic variation type estimation of $\sigma$ by employing the process $\widetilde{X}$. Indeed, by setting $f_{\theta}\equiv 1$, Theorem 6.10 immediately yields the following convergence in probability:
$\frac{3}{2}\frac{1}{N}Q_{n}^{N,P}(1)\xrightarrow{\mathbb{P}}\sigma_{0}^{2}.$
We now turn to preparing the proof of asymptotic normality, which relies on the following convergence in distribution. We begin by analyzing $I_{n}^{N,C}(f_{\theta})$ and $I_{n}^{N,P}(f_{\theta})$.
@@
$\frac{1}{\sqrt{N}}I_{n}^{N,C}(f)\xrightarrow{d}\mathcal{N}(0,\bar{\Pi}((fa_{\sigma_{0}})^{2})),\qquad\frac{1}{\sqrt{N}}I_{n}^{N,P}(f)\xrightarrow{d}\mathcal{N}(0,\bar{\Pi}((fa_{\sigma_{0}})^{2})).$
@@
$\frac{1}{\sqrt{N\Delta_{n}}}(Q_{n}^{N,C}(f)-\Delta_{n}\nu_{n}^{N,C}(fa_{\sigma_{0}}^{2}))\xrightarrow{d}\mathcal{N}(0,2\bar{\Pi}(f^{2}a^{4}_{\sigma_{0}})),$
$\frac{1}{\sqrt{N\Delta_{n}}}(Q_{n}^{N,P}(f)-\frac{2}{3}\Delta_{n}\nu_{n}^{N,P}(fa_{\sigma_{0}}^{2}))\xrightarrow{d}\mathcal{N}(0,\bar{\Pi}(f^{2}a^{4}_{\sigma_{0}})).$
@@
7 Proof of main results
@@
$\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta_{n}}}{N}L_{n}^{N,P}(\mu,\sigma)$ $\xrightarrow{\mathbb{P}}\bar{\Pi}\Big(\frac{a^{2}_{\sigma_{0}}}{a^{2}_{\sigma}}+\log a^{2}_{\sigma}\Big).$ (25)
@@
uniformly in $\theta$ according to Proposition 6.8 and Theorems 6.9 and 6.10. Similarly, in the partial observation case we have that
@@
uniformly in $(\mu,\sigma)$ according to the same Proposition 6.8 and Theorems 6.9 and 6.10.
@@
In what follows, we present the proofs of our main results for the case of scalar parameters in order to simplify the notation. For the same reason, we will omit in $\Delta_{n}$ the dependence in $n$. The scheme of the proof is the same in both the complete and the partial observation case. For $A\in\{P,C\}$, we use Taylor’s formula and, as $\nabla_{\theta}L_{n}^{N,A}(\hat{\theta}_{n}^{N,A})=0$, we obtain
@@
$\Sigma_{n}^{N,A}(\theta)$ $=M_{n}^{N,A}\nabla^{2}_{\theta}L_{n}^{N,A}(\theta)M_{n}^{N,A}$
@@
The asymptotic behavior of the estimator is given by the asymptotic behavior of the first and second derivatives of the contrast function. Let us now detail the cases of complete and partial observation separately. We start by considering the complete observation framework. Note that
@@
$-\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,C}(\theta_{0})$ $=\frac{1}{\sqrt{N\Delta}}Q_{n}^{N,C}\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{4}_{\sigma_{0}}}\Big)-\sqrt{\frac{\Delta}{N}}\nu_{n}^{N,C}\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{2}_{\sigma_{0}}}\Big)+o_{\mathbb{P}}(1).$
In the equality above, $o_{\mathbb{P}}(1)$ denotes the terms
$2\sqrt{\frac{\Delta}{N}}I_{n}^{N,C}\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}b_{\mu_{0}}}{a^{4}_{\sigma_{0}}}\Big)-\Delta\sqrt{\frac{\Delta}{N}}\nu_{n}^{N,C}\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}b_{\mu_{0}}^{2}}{a^{4}_{\sigma_{0}}}\Big),$
both of which vanish in probability as $n,N\to\infty$. This negligibility is a direct consequence of the assumption $N\Delta\to 0$, as well as the convergence results established in Theorem 6.9 and Proposition 6.8, respectively. Then, Theorem 6.11 ensures that
$\frac{1}{\sqrt{N}}\partial_{\mu}L_{n}^{N,C}(\theta_{0})\xrightarrow{d}\mathcal{N}\Big(0,\bar{\Pi}\Big(\Big(2\frac{\partial_{\mu}b_{\mu_{0}}}{(a^{2}_{\sigma_{0}})}a_{\sigma_{0}}\Big)^{2}\Big)\Big)=N\Big(0,4\bar{\Pi}\Big(\Big(\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}}\Big)^{2}\Big)\Big).$
@@
$\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,C}(\theta_{0})\xrightarrow{d}\mathcal{N}\Big(0,2\bar{\Pi}\Big(\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{2}_{\sigma_{0}}}\Big)^{2}\Big)\Big).$
@@
$\nabla_{\theta}L_{n}^{N,C}(\theta_{0})M_{n}^{N,C}\xrightarrow{d}\mathcal{N}(0,\Sigma^{C}(\theta_{0})),$
@@
$\Sigma^{C}(\theta_{0})=\begin{pmatrix}4\bar{\Pi}\big(\big(\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}}\big)^{2}\big)&0\\ 0&2\bar{\Pi}\big(\big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{2}_{\sigma_{0}}}\big)^{2}\big)\end{pmatrix}$.
@@
$\frac{1}{N}\partial^{2}_{\mu}L_{n}^{N,C}(\theta)$ $=-\frac{2}{N}I_{n}^{N,C}\Big(\frac{\partial_{\mu}^{2}b_{\mu}}{a_{\sigma}^{2}}\Big)+\frac{\Delta}{N}\nu_{n}^{N,C}\Big(\partial_{\mu}\big(2\frac{\partial_{\mu}b_{\mu}}{a_{\sigma}^{2}}(b_{\mu}-b_{\mu_{0}})\big)\Big)$
$=o_{\mathbb{P}}(1)+\frac{\Delta}{N}\nu_{n}^{N,C}\Big(\frac{2}{a_{\sigma}^{2}}\big(\partial_{\mu}^{2}b_{\mu}(b_{\mu}-b_{\mu_{0}})+(\partial_{\mu}b_{\mu})^{2}\big)\Big)$
$\xrightarrow{\mathbb{P}}\bar{\Pi}\Big(\frac{2}{a_{\sigma}^{2}}\big(\partial_{\mu}^{2}b_{\mu}(b_{\mu}-b_{\mu_{0}})+(\partial_{\mu}b_{\mu})^{2}\big)\Big)=:2\Sigma^{(1)}(\theta),$
where the $o_{\mathbb{P}}(1)$ term arises because $\frac{2}{N}I_{n}^{N,C}(\frac{\partial_{\mu}^{2}b_{\mu}}{a_{\sigma}^{2}})$ converges to zero in probability, an asymptotic property guaranteed by Theorem 6.9. The convergence holds true uniformly in $\theta$ thanks to Proposition 6.8. Observe in particular that $\Sigma^{(1)}(\theta_{0})=\bar{\Pi}(\frac{(\partial_{\mu}b_{\mu_{0}})^{2}}{a^{2}_{\sigma_{0}}})$. Furthermore,
@@
$=2\frac{\sqrt{\Delta}}{N}I_{n}^{N,C}\Big(\frac{\partial_{\mu}b_{\mu}}{a_{\sigma}^{4}}\partial_{\sigma}a_{\sigma}^{2}\Big)-\frac{\Delta^{\frac{3}{2}}}{N}\nu_{n}^{N,C}\Big(2\frac{\partial_{\sigma}a_{\sigma}^{2}\partial_{\mu}b_{\mu}}{a_{\sigma}^{4}}(b_{\mu}-b_{\mu_{0}})\Big)\xrightarrow{\mathbb{P}}0$
@@
Regarding the second derivatives in $\sigma$,
$\frac{\Delta}{N}\partial^{2}_{\sigma}L_{n}^{N,C}(\theta)=\frac{1}{N}Q_{n}^{N,C}\Big(2\frac{(\partial_{\sigma}a_{\sigma}^{2})^{2}}{a_{\sigma}^{6}}-\frac{\partial^{2}_{\sigma}a_{\sigma}^{2}}{a^{4}_{\sigma}}\Big)+o_{\mathbb{P}}(1)+\frac{\Delta}{N}\nu_{n}^{N,C}\Big(\frac{\partial^{2}_{\sigma}a_{\sigma}^{2}}{a_{\sigma}^{2}}-\frac{(\partial_{\sigma}a_{\sigma}^{2})^{2}}{a_{\sigma}^{4}}\Big),$
where we have replaced
$\frac{2\Delta}{N}I_{n}^{N,C}\Big(\frac{b_{\mu}\partial^{2}_{\sigma}a^{2}_{\sigma}}{a^{4}}-2\frac{b_{\mu}(\partial_{\sigma}a^{2}_{\sigma})^{2}}{a^{6}_{\sigma}}\Big)-\frac{\Delta^{2}}{N}\nu_{n}^{N,C}\Big(\frac{b^{2}_{\mu}\partial^{2}_{\sigma}a^{2}_{\sigma}}{a^{4}_{\sigma}}-2\frac{b^{2}_{\mu}(\partial_{\sigma}a^{2}_{\sigma})^{2}}{a^{6}_{\sigma}}\Big)$
with $o_{\mathbb{P}}(1)$ because of the convergences in probability gathered in Theorem 6.9 and Proposition 6.8, respectively. It follows that $\frac{\Delta}{N}\partial^{2}_{\sigma}L_{n}^{N,C}(\theta)$ converges in probability to
$\bar{\Pi}\Big(a^{2}_{\sigma_{0}}\Big(2\frac{(\partial_{\sigma}a_{\sigma}^{2})^{2}}{a_{\sigma}^{6}}-\frac{\partial^{2}_{\sigma}a_{\sigma}^{2}}{a^{4}_{\sigma}}\Big)\Big)+\bar{\Pi}\Big(\frac{\partial^{2}_{\sigma}a_{\sigma}^{2}}{a_{\sigma}^{2}}-\frac{(\partial_{\sigma}a_{\sigma}^{2})^{2}}{a_{\sigma}^{4}}\Big)=:\Sigma^{(2)}(\theta)$
uniformly in $\theta$, thanks to Proposition 6.8 and Theorem 6.10. We remark that $\Sigma^{(2)}(\theta)=\bar{\Pi}(\frac{(\partial_{\sigma}a^{2}_{\sigma})^{2}}{a_{\sigma}^{4}})$. This implies that the matrix $\Sigma_{n}^{N,C}(\theta)$ converges in probability, uniformly in $\theta$, towards
$\Sigma(\theta)=\begin{pmatrix}2\Sigma^{(1)}(\theta)&0\\ 0&\Sigma^{(2)}(\theta)\end{pmatrix}$.
@@
$(\Sigma(\theta_{0}))^{-1}\times\begin{pmatrix}4\Sigma^{(1)}(\theta_{0})&0\\ 0&2\Sigma^{(2)}(\theta_{0})\end{pmatrix}\times(\Sigma(\theta_{0}))^{-1}=\begin{pmatrix}(\Sigma^{(1)}(\theta_{0}))^{-1}&0\\ 0&2(\Sigma^{(2)}(\theta_{0}))^{-1}\end{pmatrix}$,
@@
$L_{n}^{N,P}(\theta)=\frac{3}{2}\frac{1}{\Delta}Q_{n}^{N,P}\Big(\frac{1}{a_{\sigma}^{2}}\Big)-3I_{n}^{N,P}\Big(\frac{b_{\mu}}{a_{\sigma}^{2}}\Big)+\frac{3}{2}\Delta\nu_{n}^{N,P}\Big(\frac{b_{\mu}}{a_{\sigma}^{2}}(b_{\mu}-2b_{\mu_{0}})\Big)+\nu_{n}^{N,P}\big(\log(a_{\sigma}^{2})\big).$
@@
$-\frac{1}{\sqrt{N}}\partial_{\mu}L_{n}^{N,P}(\theta_{0})$ $=\frac{3}{\sqrt{N}}I_{n}^{N,P}\Big(\frac{\partial_{\mu}b_{\mu_{0}}}{a^{2}_{\sigma_{0}}}\Big)\xrightarrow{d}\mathcal{N}\Big(0,9\bar{\Pi}\Big(\Big(\frac{\partial_{\mu}b_{\mu_{0}}}{a_{\sigma_{0}}}\Big)^{2}\Big)\Big).$
@@
$\sqrt{\frac{\Delta}{N}}\partial_{\sigma}L_{n}^{N,P}(\theta_{0})\xrightarrow{d}\mathcal{N}\Big(0,\frac{9}{4}\bar{\Pi}\Big(\Big(\frac{\partial_{\sigma}a^{2}_{\sigma_{0}}}{a^{2}_{\sigma_{0}}}\Big)^{2}\Big)\Big).$
Recall that Theorem 6.10 applies with limiting factor $\frac{2}{3}$ in the partial observation case. Then, acting as above, it is easy to check that, uniformly in $\theta$,
$\Sigma_{n}^{N,P}(\theta)\xrightarrow{\mathbb{P}}\begin{pmatrix}3\Sigma^{(1)}(\theta)&0\\ 0&\Sigma^{(2)}(\theta)\end{pmatrix}$
with $\Sigma^{(j)}(\theta)$ defined as in the case of complete observation we dealt with before. We conclude again by applying the inverse of $\int_{0}^{1}\Sigma_{n}^{N,P}(\theta_{0}+u(\hat{\theta}_{n}^{N,P}-\theta_{0}))du$ in Equation (30) and using the continuous mapping theorem. The result is therefore proven.
Appendix A Proof of technical results